# library.bib

### v2

@@ Abstract
We extend Friedman’s theorem to show that, for any fixed $r$, the random $2r$–regular Schreier graphs depicting the action of $r$ random permutations of $[n]$ on $k_{n}$–tuples of distinct elements in $[n]$ are a.a.s. weakly Ramanujan, for any $k_{n}\leq n^{\frac{1}{12}-\epsilon}.$ Previously this was known only for $k$–tuples where $k$ is fixed. In fact, we prove the stronger result of strong convergence of random matrices in irreducible representations of $S_{n}$ of quasi–exponential dimension.
We also give a new bound for the expected stable irreducible character of a random permutation obtained via a word map, showing that $\mathbb{E}\left[\chi^{\mu}\left(w(\sigma_{1},\dots,\sigma_{r})\right)\right]=O\left(n^{-k}\right)$, where $k$ is the number of boxes outside the first row of the Young diagram $\mu,$ solving one aspect of a conjecture of Hanany and Puder. We obtain this bound using an extension of Wise’s $w$–cycle conjecture.
@@ 1 Introduction
Let $G_{n}$ be a connected $2r$–regular graph on $n$ vertices, $A_{G_{n}}$ its adjacency matrix with eigenvalues $2r=\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{n}$. Let $\lambda(G_{n})=\mathrm{max}\left(\lambda_{2},|\lambda_{n}|\right).$ The Alon–Boppana bound dictates that
$\lambda\left(G_{n}\right)\geq 2\sqrt{2r-1}-o(1),$
so that $2r$–regular graphs with $\lambda\left(G_{n}\right)\leq 2\sqrt{2r-1}$ have optimal spectral gap, see [Alon86, Friedman03, Nilli91] for example.
Alon conjectured in [Alon86] that most $d$–regular graphs on a large number of vertices are weakly Ramanujan with high probability, and a breakthrough paper by Friedman [Friedman2008] proved Alon’s conjecture, proving that random $2r$–regular Schreier graphs (see §2.8), $\mathrm{Sch}\left(S_{n}\curvearrowright[n],\sigma_{1},\dots,\sigma_{r}\right)$ are asymptotically almost surely (a.a.s.) weakly Ramanujan, where a.a.s. means with probability $\to 1$ as $n\to\infty.$
Questions remain about the expansion properties of Cayley graphs of $S_{n}.$ These questions are very different in the sense that the Cayley graphs have $n!\sim n^{n}$ vertices, rather than only $n$ vertices. Kassabov, in [Kassabov], proved the following.
@@
Whether or not random Cayley graphs of $S_{n}$ (ie with a random generating set) are uniform expanders, or if they satisfy the conclusion of Friedman’s theorem, are still open questions. Our main result below is in this direction.
For any integers $r,k>0,$ define $\mathcal{G}_{r}\left(n,k\right)$ to be the collection of Schreier graphs $\mathrm{Sch}\left(S_{n}\curvearrowright[n]_{k},\sigma_{1},\dots,\sigma_{r}\right),$ where $[n]_{k}$ denotes the set of all $k$–tuples of distinct elements in $[n]$.
@@
Fix any integer $r>1$ and let $\alpha<\frac{1}{12}.$ For any sequence of positive integers $\left(k_{n}\right)_{n\geq 1}$ with $k_{n}\leq n^{\alpha}$, let $\left(G_{n}\right)_{n\geq 1}$ be a sequence of random $2r$–regular Schreier graphs, where for each $n\geq 1$, $G_{n}\in\mathcal{G}_{r}\left(n,k_{n}\right)$ is obtained by choosing $r$ i.i.d. uniformly random permutations, $\sigma_{1},\dots,\sigma_{r}\in S_{n}.$
@@
$\mathbb{P}\left[\lambda\left(G_{n}\right)\leq 2\sqrt{2r-1}+\epsilon\right]\overset{n\rightarrow\infty}{\longrightarrow}1.$
The case $\alpha=0$ is exactly Friedman’s theorem. The collection of Schreier graphs $\mathcal{G}_{r}\left(n,n\right)$ is the collection of Cayley graphs of $S_{n},$ so proving the above result for $\alpha=1$ would show that random Cayley graphs of $S_{n}$ are a.a.s. weakly Ramanujan. Nevertheless, Theorem 1.2 is the first result showing that the Schreier graphs constructed using random permutations acting on tuples of distinct elements have near–optimal spectral gap with high probability, in the case where the size of the tuple is allowed to grow with $n.$
Previously, this was known for tuples of size $1$ by [Friedman2008] (and later, with a shorter proof, by Bordenave [Bordenave2020]); for pairs of distinct elements by Bordenave and Collins [BordenaveCollins2018] and more recently, for $k$–tuples for any fixed $k$, by Chen, Garza–Vargas, Tropp and van Handel in [CGVTvH2024].
@@
Prior to the result for fixed $k$ in [CGVTvH2024], it was already known that for any fixed $k,$ a random regular Schreier graph of $S_{n}$ has a uniform spectral gap with high probability. Indeed, in [Friedmanetal, Theorem 2.1], they prove that, for any fixed $k,r>0,$ if $G_{n}\in\mathcal{G}_{r}\left(n,k\right)$ is a random Schreier graph, then for any $\epsilon>0,$
$\mathbb{P}\left[\lambda\left(G_{n}\right)\leq(1+\epsilon)2r\left(\frac{\sqrt{2r-1}}{r}\right)^{\frac{1}{r+1}}\right]\overset{n\to\infty}{\longrightarrow}1.$
@@
A comment on ‘derandomization’. In [Mohanty], using Bordenave’s proof of Friedman’s theorem, they give a construction of a regular graph on $O(N)$ vertices, that is near–Ramanujan with high probability (probability $\geq 1-O\left(\frac{\log N}{2^{100\left(\log N\right)^{\frac{1}{4}}}}\right)$). They obtain such a graph by randomly choosing a ‘seed’ $s\in\{0,1\}^{O\left(\log(N)^{2}\right)}$ and using this as the input to an algorithm. That is, their construction requires $\sim\log(N)^{2}$ random bits. With the construction in Theorem 1.2, one can construct a regular graph on $N$ vertices that is near–Ramanujan with high probability (probability $\geq 1-O\left(\frac{1}{\left(\log\left(N\right)^{12}\right)^{1-(12+\epsilon)\alpha}}\right)$), using $\sim\log(N)^{12}$ random bits, by sampling random permutations $\sigma_{1},\dots,\sigma_{r}\in S_{n}$, with $n\sim\log(N)^{12}$. Although this requires slightly more random bits than the algorithm given in [Mohanty], it is a more algebraic construction and illustrates a different approach to such a problem.  [^1] We thank Ramon van Handel for pointing out this application.
Our proof of Theorem 1.2 relies on the remarkable new approach to strong convergence detailed in [CGVTvH2024], see §1.1, as well as the additional criterion for temperedness of arbitrary functions on finitely generated groups $\Gamma$ with a finite fixed generating set (an adaptation of the classical notion of a tempered representation) given by Magee and de la Salle in [MAgeeDelaSalle].
To use this approach effectively, we prove our other main theorem pertaining to word maps and stable irreducible characters (see §2.9), in which we prove a new bound on the expected stable irreducible character of a $w$–random permutation. Given a Young diagram $\lambda\vdash k$ and $n\geq k+\lambda_{1},$ we define the Young diagram $\lambda^{+}(n)=\left(n-k,\lambda\right)\vdash n$, see §2.1.
@@
Let $w\in F_{r}=\langle x_{1},\dots,x_{r}\rangle$ be a word in the free group on $r$ generators with $w\neq e$. Let $k\in\mathbb{Z}_{>0}$ and $\lambda\vdash k$. Then $\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]$ is a rational expression in $n$ and, if $w$ is not a proper power, then
$\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]=O\left(\frac{1}{n^{k}}\right).$
@@
That $\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]$ can be written as a rational expression in $n$ follows from (24) or indeed by combining [Nica, LinialPuder] with [HananyPuder, Proposition B.2]. In the case of $w$ being a proper power, the conclusion of Theorem 1.5 follows from [Nica] and [LinialPuder, Section 4].
@@
We prove Theorem 1.5 using ‘combinatorial integration’. We use a projection formula obtained in our previous paper [Cassidy2023] to translate the computation of $\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]$ in to combinatorial problem. We then prove an extension of Wise’s $w$–cycles conjecture (see [LouderWilton, HelferWise, Wise]), a theorem in low–dimensional topology, to obtain our final bound.
Theorem 1.5 generalizes the work of Nica [Nica] and Linial and Puder [LinialPuder], where it is established that, if #$\#\mathrm{fix}(\sigma)$ is the number of fixed points of a permutation $\sigma\in S_{n},$ then, if $w$ is not a proper power,
$\mathbb{E}_{w}\left[\#\mathrm{fix}(\sigma)-1\right]=O\left(\frac{1}{n}\right)$ (1)
Of course, #$\#\mathrm{fix}-1$ is the character of the standard representation of $S_{n},$ $\chi^{(n-1,1)}$, so that this is a special case of our theorem. In [PuderParzanchewski2012], Puder and Parzanchevski improve this, showing that
@@
where $\pi(w)$ is the primitivity rank of the word $w$ (see, for example [PuderParzanchewski2012, Definition 1.7], for the definition). Hanany and Puder [HananyPuder, Theorem 1.3] generalize this result to other stable irreducible representations of $S_{n}.$ They show that, if $w$ is not a proper power, then for any $k\neq 0,1$ and for any $\lambda\vdash k,$
@@
Theorem 1.5 is thus an improvement on this bound in the regime $k>\pi(w).$
In [HananyPuder, Conjecture 1.8], it is conjectured that
$\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]=O\left(\frac{1}{n^{k(\pi(w)-1)}}\right).$
So Theorem 1.5 confirms the conjecture for words $w$ with primitivity rank $\pi(w)=2.$ In cases where there is no further information on $w$ other than $w\neq e$ and $w$ is not a proper power, the bound given in Theorem 1.5 is expected to be sharp. In fact, it has recently been conjectured by Puder and Shomroni (see [PuderShomroni, Conjecture 1.2]) that we should have
@@
where $\mathrm{s}\pi(w)$ is a property of $w$ known as the stable primitvity rank (this was first introduced by Wilton in [Wilton], indeed see Definition 10.6 in (ibid.) or [PuderShomroni, Definition 4.1] for an equivalent definition). They also conjecture that this bound is the best one possible. That is, for every word $w\in F_{r}$, for every $k,$ there is a Young diagram $\lambda\vdash k$ such that $\frac{1}{n^{k(\mathrm{s}\pi(w))}}=O\left(\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]\right).$ An additional conjecture of Wilton (see eg [PuderShomroni, Conjecture 4.7]) posits that, for any $w\in F_{r},$ we have the equality $\mathrm{s}\pi(w)=\pi(w)-1$, so that the conjecture of Hanany and Puder is a combination of these two conjectures.
@@ 1.1 Strong Convergence
As we noted earlier, we actually prove a stronger statement than Theorem 1.2, that of strong convergence of random representations of $F_{r}=\left\langle x_{1},\dots,x_{r}\right\rangle$ that factor through permutations.
@@
Given a sequence of random, finite dimensional, unitary representations $\{\pi_{n}:F_{r}\to U\left(N_{n}\right)\}_{n\geq 1},$ we say that $\pi_{n}$ strongly converge to the left regular representation $\lambda:F_{r}\to U\left(l^{2}\left(F_{r}\right)\right)$ a.a.s. if, for any $z\in\mathbb{C}\left[F_{r}\right]$, for any $\epsilon>0$, we have
$\mathbb{P}\left[\left|\left\|\pi_{n}(z)\right\|-\left\|\lambda(z)\right\|\right|>\epsilon\right]\overset{n\to\infty}{\longrightarrow}1.$ (2)
@@
We are interested in the cases whereby the representations $\pi_{n}=\rho\circ\phi_{n},$ where $\phi_{n}\in\mathrm{Hom}\left(F_{r},S_{n}\right)$ is random and $\rho:S_{n}\to\mathrm{End}\left(V\right)$ is a representation of $S_{n}$ of dimension $N_{n}.$ The power of strong convergence in this setting is that, since the convergence must hold for every $z\in\mathbb{C}\left[F_{r}\right]$, it allows us to prove results like Theorem 1.2, which require (2) to hold only for specific elements of $\mathbb{C}\left[F_{r}\right].$
For example, proving Friedman’s Theorem is equivalent to showing that
$\mathbb{P}\left[\left|\left\|\pi_{n}(z)\right\|-2\sqrt{2r-1}\right|>\epsilon\right]\overset{n\to\infty}{\longrightarrow}1$ (3)
@@
and $\pi_{n}=\rho^{(n-1,1)}\circ\phi_{n}$, with $\phi_{n}\in\mathrm{Hom}\left(F_{r},S_{n}\right)$ uniformly random (ie obtained by uniformly randomly choosing a permutation $\sigma_{i}\in S_{n}$ for each $i$ and mapping $x_{i}\mapsto\sigma_{i}$) and where
@@
is the standard representation of $S_{n}.$ This is because, if $\rho:S_{n}\to\mathrm{End}\left(\mathbb{C}^{n}\right)$ is the permutation representation of $S_{n},$ then
@@
the $k^{\mathrm{th}}$ tensor power of the permutation representation. We then define
$\rho_{n,k}:\mathbb{C}\left[S_{n}\right]\to\mathrm{End}\left(V_{n,k}\right)$
to be the restriction of $\bar{\rho}_{n,k}$ to the orthocomplement to the $S_{n}$–invariant vectors, $V_{n,k}\subseteq\left(\mathbb{C}^{n}\right)^{\otimes k}$ and define a random sequence of unitary representations of $F_{r},$$\{\pi_{n,k}\overset{\mathrm{def}}{=}\rho_{n,k}\circ\phi_{n}\}_{n\geq 1},$ with $\phi_{n}\in\mathrm{Hom}\left(F_{r},S_{n}\right)$ uniformly random as before.
Bordenave and Collins [BordenaveCollins2018] proved strong convergence to the left regular representation $\lambda:F_{r}\to U\left(l^{2}\left(F_{r}\right)\right)$ a.a.s. for the sequence of random representations of $F_{r}$, $\{\pi_{n,1}\}_{n\geq 1}$. They extend their results to show strong convergence a.a.s. for $\{\pi_{n,2}\}_{n\geq 1}$. With the new approach in [CGVTvH2024], strong convergence a.a.s. for $\{\pi_{n,k}\}_{n\geq 1}$ for any fixed $k$ is now known. We adopt and extend this new approach to prove our final main theorem below.
@@
For any $\alpha<\frac{1}{12}$, for any $z\in\mathbb{C}\left[F_{r}\right]$ and for any $\epsilon>0,$
$\mathbb{P}\left[\sup_{k_{n}\leq n^{\alpha}}\left|\left\|\pi_{n,k_{n}}(z)\right\|-\left\|\lambda(z)\right\|\right|<\epsilon\right]\overset{n\to\infty}{\longrightarrow}1.$

### v3

# Random permutations acting on kk–tuples have near–optimal spectral gap for $k=\mathrm{poly}(n)$

@@ Abstract
We extend Friedman’s theorem to show that, for any fixed $r>1$, a random 2r2r–regular Schreier graph associated with the action of $r$ uniformly random permutations of $[n]$ on $k_{n}$–tuples of distinct elements in $[n]$ has a near–optimal spectral gap with high probability, provided $k_{n}\leq n^{\frac{1}{20}-\epsilon}.$ Previously this was known only for kk–tuples where $k$ is fixed. In fact, we prove the stronger result of strong convergence of random permutations in irreducible representations of quasi–exponential dimension.
Along the way, we give a new bound for the expected stable irreducible character of a random permutation obtained via a word map, showing that $\mathbb{E}\left[\chi^{\mu}\left(w(\sigma_{1},\dots,\sigma_{r})\right)\right]=O\left(\frac{1}{\dim\chi^{\mu}}\right)=O\left(n^{-k}\right)$, where $k$ is the number of boxes outside the first row of the Young diagram $\mu,$ solving one aspect of a conjecture of Hanany and Puder. We obtain this bound using an extension of Wise’s ww–cycle conjecture.
@@ 1 Introduction
1.1 Random regular graphs and word maps
Let $G_{n,r}$ be a connected 2r2r–regular graph on $n$ vertices, $A_{G_{n,r}}$ its adjacency matrix with eigenvalues $2r=\lambda_{1}>\lambda_{2}\geq\dots\geq\lambda_{n}\geq-2r$. Denote $\lambda(G_{n,r})=\mathrm{max}_{|\lambda_{i}|\neq 2r}|\lambda_{i}|$ the largest non–trivial eigenvalue. The Alon–Boppana bound dictates that
$\lambda\left(G_{n,r}\right)\geq 2\sqrt{2r-1}-o(1),$
so that 2r2r–regular graphs with $\lambda\left(G_{n,r}\right)\leq 2\sqrt{2r-1}$ have optimal spectral gaps, see [Alo86, Fri03, Nil91] for example.
In [Alo86], Alon conjectured that a random 2r2r–regular graph on $n$ vertices, $G_{n,r},$ has a near–optimal spectral gap (i.e. $\lambda\left(G_{n,r}\right)\leq 2\sqrt{2r-1}+\epsilon$) with high probability (meaning with probability $\to 1$ as $n\to\infty$). Alon’s conjecture was resolved by Friedman [Fri08].
Theorem 1.1 (Friedman).
If $G_{n,r}$ is a random 2r2r–regular graph on $n$ vertices, then for any $\epsilon>0,$
$\mathbb{P}\left[\lambda(G_{n,r})\leq 2\sqrt{2r-1}+\epsilon\right]\overset{n\to\infty}{\longrightarrow}1.$
To sample a random 2r2r–regular graph Friedman uses the permutation model, taking $G_{n,r}$ to be a random Schreier graph associated to the action of $r$ uniformly random permutations on $[n]$, denoted $\mathrm{Sch}\left(S_{n}\curvearrowright[n],\sigma_{1},\dots,\sigma_{r}\right)$, see §2.8. Questions remain about the expansion properties of Cayley graphs of $S_{n}.$ These questions are very different in the sense that the Cayley graphs have $n!$ vertices rather than $n$, and to answer them, one must have some control over all non–trivial finite dimensional complex representations of $S_{n}.$ Kassabov, in [Kas05], proved that explicit generating sets of a bounded size can be constructed such that the corresponding Cayley graphs of $S_{n}$ form a family of bounded–degree expanders.
@@ 1.1 Random regular graphs and word maps
Whether or not this is typical behaviour, that is, whether or not Cayley graphs of $S_{n}$ with a random generating set of fixed size have a uniform spectral gap with high probability, remains an open problem. Our main result below is in this direction. For any integers $r,k>0,$ define $\mathcal{G}_{r}\left(n,k\right)$ to be the collection of Schreier graphs $\mathrm{Sch}\left(S_{n}\curvearrowright[n]_{k},\sigma_{1},\dots,\sigma_{r}\right),$ where $n]_{k}$ denotes the set of all kk–tuples of distinct elements in $[n]$ and $\sigma\left(i_{1},\dots,i_{k}\right)=\left(\sigma(i_{1}),\dots,\sigma(i_{k})\right).$
@@
Fix any integer $r>1$ and let $\alpha<\frac{1}{20}.$ For any sequence of positive integers $\left(k_{n}\right)_{n\geq 1}$ with $1\leq k_{n}\leq n^{\alpha}$, let $\left(G_{n,r,k_{n}}\right)_{n\geq 1}$ be a sequence of random 2r2r–regular Schreier graphs, where for each $n\geq 1$, $G_{n,r,k_{n}}\in\mathcal{G}_{r}\left(n,k_{n}\right)$ is obtained by choosing $r$ i.i.d. uniformly random permutations, $\sigma_{1},\dots,\sigma_{r}\in S_{n}.$
@@
$\mathbb{P}\left[\lambda\left(G_{n,r,k_{n}}\right)\leq 2\sqrt{2r-1}+\epsilon\right]\overset{n\rightarrow\infty}{\longrightarrow}1.$
One sees that $n]_{n}=S_{n}$ and so, in this case, a randomly sampled Schreier graph is simply a random Cayley graph of $S_{n}$ of degree $2r$. Thus, the above result for $\alpha=1$ ($k_{n}=n)$ would imply that, for each fixed $r>1,$ not only are random 2r2r–regular Cayley graphs of $S_{n}$ uniform expanders with high probability, but that they have a near–optimal spectral gap with high probability. As we mentioned above, this question is still wide open, but we include this discussion for motivation.
Previously, Theorem 1.3 was known for tuples of size $1$ (i.e. with $\alpha=0)$, which is exactly Friedman’s theorem [Fri08] (see also the shorter proof given by Bordenave [Bor20]); for pairs of distinct elements by Bordenave and Collins [BC19] and, more recently, for kk–tuples for any fixed $k$, by Chen, Garza–Vargas, Tropp and van Handel in [CGVTvH24]. One can then consider Theorem 1.3 as progress towards bridging the substantial gap between these results and the open questions regarding the spectral gap of random fixed–degree Cayley graphs of $S_{n}$. Indeed, we now know that the random regular Schreier graphs described above have a near–optimal spectral gap with high probability when the size $k$ of the tuple is allowed to grow with $n$, and we can even take $k$ to grow polynomially with $n$.
@@
Prior to the result for fixed $k$ in [CGVTvH24], it was previously known that for any fixed $k,$ a random 2r2r–regular Schreier graph of $S_{n}$ has a uniform spectral gap with high probability. Indeed, in [FJR+98, Theorem 2.1], they prove that, for any fixed $k>0$ and $r>1$, if $G_{n,r,k}\in\mathcal{G}_{r}\left(n,k\right)$ is a random 2r2r–regular Schreier graph, then for any $\epsilon>0,$
$\mathbb{P}\left[\lambda\left(G_{n,r,k}\right)\leq(1+\epsilon)2r\left(\frac{\sqrt{2r-1}}{r}\right)^{\frac{1}{r+1}}\right]\overset{n\to\infty}{\longrightarrow}1.$
@@
A comment on ‘derandomization’. In [MOP20], using Bordenave’s proof of Friedman’s theorem, they give a construction of a random regular graph on $O(N)$ vertices, that has a near–optimal spectral gap with high probability (probability $\geq 1-O\left(\frac{\log N}{2^{100\left(\log N\right)^{\frac{1}{4}}}}\right)$). They obtain such a graph by randomly choosing a ‘seed’ $s\in\{0,1\}^{O\left(\log(N)^{2}\right)}$ and using this as the input to an algorithm. That is, their construction requires $\sim\log(N)^{2}$ random bits. With the construction in Theorem 1.3, one can construct a regular graph on $N$ vertices with a near–optimal spectral gap with high probability (probability $\geq 1-O\left(\frac{1}{\left(\log\left(N\right)^{20}\right)^{1-(20+\epsilon)\alpha}}\right)$), using $\sim\log(N)^{20}$ random bits, by sampling random permutations $\sigma_{1},\dots,\sigma_{r}\in S_{n}$, with $n\sim\log(N)^{20}$. Although this requires slightly more random bits than the algorithm given in [MOP20], it is a more algebraic construction and illustrates a different approach to such a problem.  [^1] We thank Ramon van Handel for pointing out this application.
Our proof of Theorem 1.3 relies on the remarkable new approach to strong convergence detailed in [CGVTvH24] referred to as the ‘polynomial method’, see §1.2, as well as the additional criterion for temperedness of arbitrary functions on finitely generated groups $\Gamma$ with a finite fixed generating set (an adaptation of the classical notion of a tempered representation) given by Magee and de la Salle in [MdlS24].
To use this approach effectively, we prove our other main theorem pertaining to word maps and stable irreducible characters (see §2.9), in which we prove a new bound on the expected stable irreducible character of a ww–random permutation. Given a Young diagram $\lambda\vdash k$ and $n\geq k+\lambda_{1},$ we define the Young diagram $\lambda^{+}(n)=\left(n-k,\lambda\right)\vdash n$, see §2.1.
@@
Let $w\in F_{r}=\langle x_{1},\dots,x_{r}\rangle$ be a word in the free group on $r$ generators with $w\neq e$. Fix $k\in\mathbb{Z}_{>0}$ and $\lambda\vdash k$. If $w$ is not a proper power, then
$\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]\overset{\mathrm{def}}{=}\mathop{\mathop{\mathbb{E}}_{\sigma_{1},\dots,\sigma_{r}\in S_{n}}\left[\chi^{\lambda^{+}(n)}\left(w\left(\sigma_{1},\dots,\sigma_{m}\right)\right)\right]}=\left(\frac{1}{\dim\chi^{\lambda^{+}(n)}}\right)=O\left(\frac{1}{n^{k}}\right).$
@@
That $\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]$ can be written as a rational expression in $n$ follows from (23) or indeed by combining [Nic94, LP10] with [HP23, Proposition B.2]. In the case of $w$ being a proper power, the conclusion of Theorem 1.6 follows from [Nic94] and [LP10, Section 4].
@@
We prove Theorem 1.6 using ‘combinatorial integration’. We use a projection formula obtained in our previous paper [Cas25] to translate the computation of $\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]$ in to combinatorial problem. We then prove an extension of Wise’s ww–cycles conjecture (see [LW17, HW16, Wis05]), a theorem in low–dimensional topology, to obtain our final bound.
Theorem 1.6 generalizes the work of Nica [Nic94] and Linial and Puder [LP10], where it is established that, if $\mathrm{fix}(\sigma)$ is the number of fixed points of a permutation $\sigma\in S_{n},$ then, if $w$ is not a proper power,
$\mathbb{E}_{w}\left[\#\mathrm{fix}(\sigma)-1\right]=O\left(\frac{1}{n}\right).$ (1)
Of course, $\mathrm{fix}-1$ is the character of the (n1)(n-1)–dimensional standard representation of $S_{n},$ $\chi^{(n-1,1)}$, so that this is a special case of our Theorem 1.6.222It is exactly this result of Nica that is used in [CGVTvH24] to give a substantially shorter proof of Friedman’s theorem. In [PP15], Puder and Parzanchevski prove that this decay is actually much faster depending on an algebraic property of the fixed word $w.$ They show that
@@
where $\pi(w)$ is the primitivity rank of the word $w$ (see, for example [PP15, Definition 1.7], for the definition). Hanany and Puder [HP23, Theorem 1.3] generalize this result to other stable irreducible representations of $S_{n}.$ They show that, if $w$ is not a proper power, then for any $k\neq 0,1$ and for any $\lambda\vdash k,$
@@
Theorem 1.6 is thus an improvement on this bound in the regime $k>\pi(w).$ Perhaps more importantly though, it gives a bound that improves when considering stable irreducible representations of higher dimensions.
In [HP23, Conjecture 1.8], it is conjectured that
$\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]=O\left(\frac{1}{\left(\dim\chi^{\lambda^{+}(n)}\right)^{\pi(w)-1}}\right)=O\left(\frac{1}{n^{k(\pi(w)-1)}}\right).$
So Theorem 1.6 confirms the conjecture for words $w$ with primitivity rank $\pi(w)=2.$ In cases where there is no further information on $w$ other than $w\neq e$ and $w$ is not a proper power, the bound given in Theorem 1.6 is expected to be sharp. In fact, it has recently been conjectured by Puder and Shomroni (see [PS23, Conjecture 1.2]) that we should have
@@
where $\mathrm{s}\pi(w)$ is a property of $w$ known as the stable primitvity rank (this was first introduced by Wilton in [Wil24], indeed see Definition 10.6 in (ibid.) or [PS23, Definition 4.1] for an equivalent definition). They also conjecture that this bound is the best one possible. That is, for every word $w\in F_{r}$, for every $k,$ there is a Young diagram $\lambda\vdash k$ such that $\frac{1}{n^{k(\mathrm{s}\pi(w))}}=O\left(\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]\right).$ An additional conjecture of Wilton (see e.g. [PS23, Conjecture 4.7]) posits that, for any $w\in F_{r},$ we have the equality $\mathrm{s}\pi(w)=\pi(w)-1$, so that the conjecture of Hanany and Puder is a combination of these two conjectures.
@@ 1.2 Strong Convergence
As we noted earlier, we actually prove a stronger statement than Theorem 1.3, that of strong convergence of random permutation representations of $F_{r}=\left\langle x_{1},\dots,x_{r}\right\rangle$.
@@
Given a sequence of random, finite dimensional, unitary representations $\{\pi_{n}:F_{r}\to U\left(N_{n}\right)\}_{n\geq 1},$ we say that $\pi_{n}$ strongly converge to the left regular representation $\lambda:F_{r}\to U\left(\ell^{2}\left(F_{r}\right)\right)$ asymptotically almost surely (a.a.s.) if, for any $z\in\mathbb{C}\left[F_{r}\right]$, for any $\epsilon>0$, we have
$\mathbb{P}\left[\big|\left\|\pi_{n}(z)\right\|-\left\|\lambda(z)\right\|\big|<\epsilon\right]\overset{n\to\infty}{\longrightarrow}1.$ (2)
@@
We are interested in the cases whereby the representations $\pi_{n}$ factor as $\pi_{n}=\rho\circ\phi_{n},$ where $\phi_{n}\in\hom\left(F_{r},S_{n}\right)$ is random and $\rho:S_{n}\to\mathrm{End}\left(V\right)$ is a representation of $S_{n}$ of dimension $N_{n}.$ The power of strong convergence in this setting is that, since the convergence must hold for every $z\in\mathbb{C}\left[F_{r}\right]$, it allows us to prove results like Theorem 1.3, which require (2) to hold only for specific elements of $\mathbb{C}\left[F_{r}\right].$
For example, Friedman’s Theorem combined with the Alon–Boppana bound can be written
$\mathbb{P}\left[\big|\left\|\pi_{n}(z)\right\|-2\sqrt{2r-1}\big|<\epsilon\right]\overset{n\to\infty}{\longrightarrow}1$ (3)
@@
and $\pi_{n}=\rho^{(n-1,1)}\circ\phi_{n}$, with $\phi_{n}\in\hom\left(F_{r},S_{n}\right)$ uniformly random (i.e. obtained by uniformly randomly choosing a permutation $\sigma_{i}\in S_{n}$ for each $i$ and mapping $x_{i}\mapsto\sigma_{i}$) and where
@@
is the standard representation of $S_{n}.$ This is because, if $\rho:S_{n}\to\mathrm{End}\left(\mathbb{C}^{n}\right)$ is the defining representation of $S_{n},$ then
@@
the $k^{\mathrm{th}}$ tensor power of the defining representation. We then define
$\rho_{n,k}:\mathbb{C}\left[S_{n}\right]\to\mathrm{End}\left(\left(\mathbb{C}^{n}\right)^{\otimes k}_{\perp\mathrm{triv}}\right)$
to be the restriction of $\bar{\rho}_{n,k}$ to the orthogonal complement to the $S_{n}$–invariant vectors, and define a random sequence of unitary representations of $F_{r},$ $\{\pi_{n,k}\overset{\mathrm{def}}{=}\rho_{n,k}\circ\phi_{n}\}_{n\geq 1},$ with $\phi_{n}\in\hom\left(F_{r},S_{n}\right)$ uniformly random as before.
In this language, Bordenave and Collins [BC19] proved strong convergence to the left regular representation $\lambda:F_{r}\to U\left(\ell^{2}\left(F_{r}\right)\right)$ a.a.s. for the sequence of random representations of $F_{r}$, $\{\pi_{n,1}\}_{n\geq 1}$. They extend their results to show strong convergence a.a.s. for $\{\pi_{n,2}\}_{n\geq 1}$. With the new ‘polynomial method’ developed in [CGVTvH24], strong convergence a.a.s. for $\{\pi_{n,k}\}_{n\geq 1}$ for any fixed $k$ is now known. We extend their method and combine with analytic results of Magee and de la Salle [MdlS24] to prove our final main theorem below.
@@
For any $\alpha<\frac{1}{20}$, for any $z\in\mathbb{C}\left[F_{r}\right]$ and for any $\epsilon>0,$
$\mathbb{P}\left[\sup_{k_{n}\leq n^{\alpha}}\bigg|\left\|\pi_{n,k_{n}}(z)\right\|-\left\|\lambda(z)\right\|\bigg|<\epsilon\right]\overset{n\to\infty}{\longrightarrow}1.$
A simple corollary of the above theorem is that one can establish the strong convergence a.a.s. to the left regular representation of $\rho\circ\phi_{n}$, for any non–trivial irreducible representation $\rho$ of $S_{n},$ provided $\dim\rho$ grows at most quasi–exponentially in $n$. This is a vast improvement on the previous results in [CGVTvH24], whereby this fact is established for $\dim\rho$ growing at most polynomially in $n$ (for polynomials of any fixed degree).
Corollary 1.11.
For any $\alpha<\frac{1}{20}$, for any $z\in\mathbb{C}\left[F_{r}\right]$ and for any $\epsilon>0,$
$\mathbb{P}\left[\sup_{\overset{\rho}{{\scriptscriptstyle\dim\rho\leq C\exp\left(n^{\alpha}\right)}}}\bigg|\left\|\rho\circ\phi_{n}(z)\right\|-\left\|\lambda(z)\right\|\bigg|<\epsilon\right]\overset{n\to\infty}{\longrightarrow}1.$
where the supremum is over all non–trivial irreducible representations of $S_{n}$.
Proof.
See the proof in [vH25, Theorem 5.8]. ∎